# DBAW-PIKAN: Dynamic Balance Adaptive Weight Kolmogorov-Arnold Neural Network for Solving Partial Differential Equations

### v1

@@ 3.3 The Helmholtz Equation
$\nabla^{2}u+k^{2}u=q(x,y),\quad(x,y)\in\Omega=[-1,1]^{2}$ (25)
@@
$q(x,y)=-(a_{1}\pi)^{2}u-(a_{2}\pi)^{2}u+k^{2}u$ (26)
Following [raissi2019physics], the analytical solution for this equation is given by:
$u(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y)$ (27)
@@
(a) Evolution of adaptive weights
(b) Evolution of the dynamic upper bound
(c) Evolution of loss components
@@ 4 Conclusion
Addressing the dual challenges of architectural expressivity bottlenecks and multi-objective optimization imbalance encountered by Physics-Informed Neural Networks (PINNs) in solving partial differential equations (PDEs), this study presents an integrated framework termed DBAW-PIKAN (Dynamic Balancing Adaptive Weighting Physics-Informed Kolmogorov-Arnold Network). This framework synergistically combines the superior function approximation capabilities of the Kolmogorov-Arnold Network (KAN) with a novel Dynamic Balancing Adaptive Weighting (DBAW) strategy characterized by a dynamically decaying upper bound constraint.
@@
  • 1.
    Architectural Innovation: By replacing the traditional Multi-Layer Perceptron (MLP) with KAN as the backbone architecture, DBAW-PIKAN effectively captures the intricate features of PDE solutions and significantly mitigates the spectral bias issue.
  • 2.
@@
Despite the promising results, this study opens several avenues for future exploration. First, while the DBAW strategy proved robust in our experiments, it introduces hyperparameters such as $\gamma_{\max}$, $\gamma_{\min}$, and $\alpha$, the optimal selection of which currently relies on empirical tuning. Future research could investigate automated hyperparameter optimization or explore self-adaptive constraint mechanisms that eliminate the need for manual decay rate settings. Second, this work primarily validates the model on (1+1)-dimensional problems; its scalability and computational efficiency for high-dimensional PDEs warrant further evaluation. Finally, applying the DBAW-PIKAN framework to more complex real-world engineering scenarios—such as turbulence simulation, multi-physics coupling, or parameter inversion from experimental data—will be a crucial next step in verifying its practical utility in science and engineering. In summary, the DBAW-PIKAN framework provides an effective solution to the core challenges facing PINNs, demonstrating the significant potential of combining advanced network architectures with carefully designed optimization strategies and offering a new perspective for the development of scientific machine learning.

### v2

@@ 3.3 The Helmholtz Equation
$\nabla^{2}u+k^{2}u=q(x,y),\quad(x,y)\in\Omega=[-1,1]^{2},$ (25)
@@
$q(x,y)=-(a_{1}\pi)^{2}u-(a_{2}\pi)^{2}u+k^{2}u.$ (26)
Following [20], the analytical solution for this equation is given by:
$u(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y).$ (27)
@@
(a) Evolution of adaptive weights
(b) Evolution of the dynamic upper bound
(c) Evolution of loss components
@@ 4 Conclusion
Addressing the dual challenges of architectural expressivity bottlenecks and multi-objective optimization imbalance encountered by PINNs in solving PDEs, we present an integrated framework termed DBAW-PIKAN . This framework synergistically combines the superior function approximation capabilities of KAN with a novel DBAW strategy characterized by a dynamically decaying upper bound constraint.
@@
  • Architectural Innovation: By replacing the traditional MLP with KAN as the backbone architecture, DBAW-PIKAN effectively captures the intricate features of PDE solutions and significantly mitigates the spectral bias issue.
@@
Despite the promising results, we open several avenues for future exploration. First, while the DBAW strategy proved robust in our experiments, it introduces hyperparameters such as $\gamma_{\max}$, $\gamma_{\min}$, and $\alpha$, the optimal selection of which currently relies on empirical tuning. Future research could investigate automated hyperparameter optimization or explore self-adaptive constraint mechanisms that eliminate the need for manual decay rate settings. Second, this work primarily validates the model on (1+1)-dimensional problems; its scalability and computational efficiency for high-dimensional PDEs warrant further evaluation. Finally, applying the DBAW-PIKAN framework to more complex real-world engineering scenarios—such as turbulence simulation, multi-physics coupling, or parameter inversion from experimental data—will be a crucial next step in verifying its practical utility in science and engineering. In summary, the DBAW-PIKAN framework provides an effective solution to the core challenges facing PINNs, demonstrating the significant potential of combining advanced network architectures with carefully designed optimization strategies and offering a new perspective for the development of scientific machine learning.
Acknowledgements
This work is supported by the Scientific Research Foundation of Fujian University of Technology, China (Grant No. GY-Z24010). The authors also thank the anonymous reviewers for their constructive comments that helped improve this paper.